The compound in which store in shows the oxidation number of + 7?

Grade:10

1 Answers

Vikas TU
7 years ago
 compound is (NH4)2SO5
and oxidation number is:
x – 5*2 = -2
x = -2 + 10 = +8
that is not the actual oxidation state however.
After Drawing te lewis dot structure for SO5^2-
U would know that S is sharing 3 double bonds with 3 O atoms and one single bond with another O atom
Hence oxidation bcomes +7 as O is more electronegative than Sulphur.